OpenStudy (anonymous):

Simplify and find The undefined values of 4x^4+6x^3/6X^2+9x

OpenStudy (lgbasallote):

hint it would be undefined when 6x^2 + 9x = 0

OpenStudy (anonymous):

2x^3(2x+3)/3x(2x+3) =>2x^2/3 if 2x+3=0 then it would be undefined

OpenStudy (anonymous):

So the should be 2x/3 no underfined values

OpenStudy (anonymous):

or 2/3x,=0

OpenStudy (anonymous):

so wat is exactly the answer

OpenStudy (anonymous):

nope answer of simplification is 2x^2/3 answer of undefined 2x-3=0